what is the least number divisible by 2,3,4,5,6,8,9,10 and what would answer be if 7 was included please have them in separate answers
Answer by CubeyThePenguin(3113)

2 = 2
3 = 3
4 = 2^2
5 = 5
6 = 2 * 3
7 = 7
8 = 2^3
9 = 3^2
10 = 2 * 5

LCM(2, 3, 4, 5, 6, 8, 9, 10) = 2^3 * 3^2 * 5 = 360

LCM(2, 3, 4, 5, 6, 7, 8, 9, 10) = 2^3 * 3^2 * 5 * 7 = 2520